OpenStudy (anonymous):

In a group of 44 students, 26 are boys. If 33 of the students have a driver’s license and if 5 girls do not have a driver’s license, how many boys have a driver’s license?

OpenStudy (anonymous):

20 boys have a driver licence.

OpenStudy (anonymous):

No. of students who don't have DL=total students - students having DL = 44-33=11 now, in 11 students , girls=5 so boys will be 11-5=6 it means 6 boys don't have a driver licence. so, total no. of boys who have driver licence= 26-6== 20